Scores of an achievement test show that it follows a normal distribution. Its mean is 78 with a standard deviation of 8. Find the interval between the middle 80% of the scores lie.

1
Expert's answer
2021-11-03T07:35:07-0400

The middle 80% implies that we are looking for the probability for 40% of the scores either side of the mean. The z-core related to 40% is


"P(Z<z)=0.5-0.4=0.1"

"z=-1.28155"

"\dfrac{x-78}{8}=-1.28155"

"x=78-1.28155(8)"

"x=67.7476"


"\dfrac{x-78}{8}=1.28155"

"x=78+1.28155(8)"

"x=88.2524"

The interval is "(67.7476, 88.2524)."